# Discrete Diffusion-Based Model-Level Explanation of Heterogeneous GNNs with Node Features

### v1

@@ Appendix F Validity Analysis
We observe that a higher number of graphs generated using DiGNNExplainer are valid (i.e., consistent with the metagraph of the dataset) for DBLP compared to IMDB (Table 7). For homogeneous datasets, the number of connected graphs is the same as valid graphs, as homogeneous graphs do not require to be validated using a metagraph.
Table 7: Analysis of the validity of generated graphs according to the metagraph.
@@
Appendix G Case Study
Example Explanations with Node Features
In this section, we visualize the feature distributions of all node types in the explanation graph that maximizes the GNN prediction for a class. We show the feature vector of the Author node with the highest probability, the most important feature in all Paper nodes, and the continuous distribution of all Term nodes.
Figures 3, 4, 5, and 6 show the explanation graphs and node features for the DiTabDDPM (50 discr. feat.) approach for DBLP. The size of the explanation graphs is between 12 and 15. All Author classes are predicted with a probability of 1.0. The ground-truth faithfulness obtained for Author classes is 0.8275. Figure 4 shows the bag-of-words features, 0.0 and 1.0, of the Author node predicted with maximum probability for each class. In Figure 5, the samples on the y-axis indicate the number of Paper nodes for a class. The most important feature of all Paper nodes has the bag-of-words features, 0.0, 2.0, 3.0, 4.0, 5.0. Class 0 has 4 Paper nodes, Class 1 has 9 Paper nodes, and Class 2 has 5 Paper nodes, and Class 3 has 6 Paper nodes. The Term nodes have continuous features (Figure 6) with a mean between 0.04 and 0.11 and a standard deviation between 0.60 and 1.22. The explanations for all approaches and all datasets can be found in our evaluation results [^7] https://github.com/ds-jrg/DiGNNExplainer/tree/main/evaluation.
Nodes 12
Prob. 1.0
(a) class 0
Nodes 15
Prob. 1.0
(b) class 1
Nodes 14
Prob. 1.0
(c) class 2
Nodes 14
Prob. 1.0
(d) class 3
Figure 3: Explanation graphs for each class of Author node of DBLP dataset that maximizes the prediction of the class. The node color blue is Paper node, red is Author node, dark green is Term node, orange is Conference node.
(a) class 0
(b) class 1
(c) class 2
(d) class 3
Figure 4: Frequency distribution plots to visualize the synthetic bag-of-words node feature vector of the Author node of the explanation graph that is classified with maximum probability for a class.
(a) class 0
(b) class 1
(c) class 2
(d) class 3
Figure 5: Frequency distribution plots to visualize the most important synthetic bag-of-words node feature vector (selected using our feature selection technique) of all Paper nodes of explanation graphs of each class. The number of samples indicate the number of Paper nodes of the explanation graph.
(a) class 0
(b) class 1
(c) class 2
(d) class 3
Figure 6: Density plots to visualize the continuous node feature vector for Term node type of the explanation graphs. We show the distribution of the Term nodes of the explanation graph that maximizes the prediction of each class.
Explanations for the same class
Table 8 shows a qualitative analysis of the explanation graphs obtained for the same class (i.e., Class 1) for all datasets and all approaches, along with example motifs for each dataset.
Table 8: Explanation graphs for Class 1
@@ Appendix G Case Study
Appendix H Runtime Analysis
Table 9 shows the breakdown of runtime for each component of our approach, and Table 10 shows runtime comparisons with baseline approaches. We would like to elaborate on the scalability of DiGNNExplainer as follows:
  • DiGNNExplainer runs the Forest Fire Sample Algorithm. The Forest Fire Algorithm is run for every node that is to be classified (e.g., author nodes in DBLP). Moreover, the Forest Fire algorithm is run for every graph size of an explanation graph (e.g., 10-15, 6 graph sizes). The runtime of the Forest Fire Sampling algorithm can be considered to be constant (we only sample small graphs of up to 15 nodes). Hence, this step is linear in the number of nodes to be classified and linear in the number of graph sizes of the explanation graphs.
  • We train DiGress on these sampled graphs. The training is linear in the number of sampled graphs and the number of steps in the diffusion process.
  • We generate explanation graphs with DiGress. The runtime is linear both in the number of generated graphs and in the number of steps in the diffusion process.
  • We train a model for node feature generation - with the runtime depending on hyperparameters and the number of node features.
  • We check whether the generated graph is consistent with the metagraph. While this is a problem of graph isomorphism checking and has an exponential runtime in the worst case, for our small explanation graphs, this can be done quickly (see table of runtimes below).
  • For each explanation graph, we make a prediction with the GNN and compute ground-truth faithfulness to select the best explanation graph.
From Table 9, we can see that the “bottleneck” is the training of the graph generation.
Table 9: Breakdown of runtime for each component of DiGNNExplainer for all datasets.
Forest Fire Graph Generation Graph Generation Node Feature Node Feature Generation Post- Graph
Sampler Training (after training) Generation Training (after training) processing selection
(secs) (mins) (mins) (secs) (secs) (secs) (secs)
DBLP 7.92 88.29 4.25 252.10 41.01 20.48 21.14
IMDB 5.08 67.58 3.2 108.73 36.46 304.57 305.07
MUTAG - 115.17 3.5 - - - 0.83
BA-Shapes 3.80 73 4.21 - - - 0.69
Tree-Cycle 3.65 82.1 4.23 - - - 0.91
Tree-Grids 3.75 87.3 4.3 - - - 0.79
BA-3Motif - 32.25 0.83 - - - 1.49
Table 10: Runtime (in secs.) comparisons for all graph sizes with baseline approaches.
XGNN 1102.26 362.58 26.56 350.82 119.04 155.4 79.47
GNNInterpreter 78.54 101.82 5.03 142.74 72.66 73.2 13.20
D4Explainer 39174.97 19406.46 658.26 5702.66 658.59 821.59 800.71
DiGNNExplainer 5895.05 5006.71 6910.2 4637.09 5191.56 5500.54 1992.29
Appendix I Ablation Study
We conduct an ablation study to investigate the effect of using multiple DiGress models for different graph sizes versus a fixed-size graph model. Moreover, we investigate the impact of using a simple GCN instead of GraphSAGE (for DBLP) and HAN (for IMDB). We show Predictive Faithfulness and Ground-truth Faithfulness scores for -
  • A fixed graph size (Table 11) : 12 for DBLP and 7 for IMDB by training a single diffusion model.
  • Multiple graph sizes (10-15)(Table 12), by training a single diffusion model. We obtain a lower ground-truth faithfulness score of 0.730 for DBLP, compared to 0.882 (Table 2 in main paper) in our main experiments, which train a separate DiGress model for each graph size (10-15).
  • Training a GCN model (Table 13) for heterogeneous datasets.
The fraction of unique and non-isomorphic graphs sampled for DBLP (graph size 12) is 0.850, and for IMDB (graph size 7) is 0.026.
Table 11: Table showing Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) for ablation study using a fixed graph size by training single Diffusion and VAE models.
@@ Appendix I Ablation Study
DiTabDDPM (50 feat.) 0.999 $\pm$ 0.0 0.717 $\pm$ 0.040
DiTabDDPM (8 discr. feat.) 0.872 $\pm$ 0.102 0.787 $\pm$ 0.037
DiTabDDPM (4 discr. feat.) 0.770 $\pm$ 0.141 0.792 $\pm$ 0.043
DiTabDDPM (2 discr. feat.) 0.925 $\pm$ 0.073 0.735 $\pm$ 0.066
VAE (baseline) 0.663 $\pm$ 0.036 0.695 $\pm$ 0.014
@@
DiTabDDPM (3066 feat., all) 0.976 $\pm$ 0.058 0.860 $\pm$ 0.051
DiTabDDPM (10 discr. feat.) 0.967 $\pm$ 0.056 0.863 $\pm$ 0.069
DiTabDDPM (5 discr. feat.) 0.951 $\pm$ 0.087 0.886 $\pm$ 0.022
DiTabDDPM (2 discr. feat.) 0.902 $\pm$ 0.099 0.836 $\pm$ 0.037
VAE (baseline) 0.999 $\pm$ 0.0 0.406 $\pm$ 0.013
Table 12: Table showing Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) for ablation study using multiple graph sizes by training a single Diffusion model.
[…]
DBLP
DiTabDDPM (50 feat.) 0.999 $\pm$ 0.0 0.730 $\pm$ 0.036
DiTabDDPM (8 discr. feat.) 0.971 $\pm$ 0.030 0.777 $\pm$ 0.043
DiTabDDPM (4 discr. feat.) 0.770 $\pm$ 0.110 0.829 $\pm$ 0.071
DiTabDDPM (2 discr. feat.) 0.938 $\pm$ 0.071 0.812 $\pm$ 0.057
IMDB
DiTabDDPM (3066 feat., all) 0.983 $\pm$ 0.021 0.853 $\pm$ 0.026
DiTabDDPM (10 discr. feat.) 0.972 $\pm$ 0.050 0.853 $\pm$ 0.030
DiTabDDPM (5 discr. feat.) 0.956 $\pm$ 0.065 0.883 $\pm$ 0.026
DiTabDDPM (2 discr. feat.) 0.934 $\pm$ 0.100 0.876 $\pm$ 0.033
Table 13: Table showing Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) for ablation study using using GCN for heterogeneous datasets in DiGNNExplainer.
[…]
DBLP
motifs per class 0.641 $\pm$ 0.022 0.8 $\pm$ 0.061
common motifs 0.632 $\pm$ 0.021 0.612 $\pm$ 0.027
IMDB
motifs per class 0.410 $\pm$ 0.010 0.850 $\pm$ 0.026
common motifs 0.400 $\pm$ 0.006 0.854 $\pm$ 0.023
Appendix J Additional Experiments
The following additional experiments compare the results for not considering and considering the dependence of node features on graph structure.
Experiments that do not consider dependence of node features
@@
Table 14 shows results for Predictive Faithfulness (PF), Ground-truth Faithfulness (GF), and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency.
Table 14: Table showing results for Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency.
[…]
DBLP
DiTabDDPM (50 feat.) 0.902 $\pm$ 0.102 0.867 $\pm$ 0.046 0.233
DiTabDDPM (8 discr. feat.) 0.797 $\pm$ 0.062 0.852 $\pm$ 0.054 0.355
DiTabDDPM (4 discr. feat.) 0.770 $\pm$ 0.121 0.820 $\pm$ 0.055 0.387
DiTabDDPM (2 discr. feat.) 0.972 $\pm$ 0.041 0.845 $\pm$ 0.052 0.384
TabDDPM (50 feat.) 0.998 $\pm$ 0.001 0.842 $\pm$ 0.054 0.194
TabDDPM (8 discr. feat.) 0.976 $\pm$ 0.034 0.852 $\pm$ 0.052 0.221
TabDDPM (4 discr. feat.) 0.916 $\pm$ 0.072 0.857 $\pm$ 0.057 0.310
TabDDPM (2 discr. feat.) 0.969 $\pm$ 0.038 0.822 $\pm$ 0.042 0.235
VAE (baseline) 0.664 $\pm$ 0.048 0.717 $\pm$ 0.044 0.0
IMDB
DiTabDDPM (3066 feat., all) 0.955 $\pm$ 0.084 0.890 $\pm$ 0.029 0.029
DiTabDDPM (10 discr. feat.) 0.977 $\pm$ 0.066 0.933 $\pm$ 0.036 0.083
DiTabDDPM (5 discr. feat.) 0.970 $\pm$ 0.065 0.900 $\pm$ 0.036 0.069
DiTabDDPM (2 discr. feat.) 0.984 $\pm$ 0.039 0.913 $\pm$ 0.047 0.040
TabDDPM (3066 feat., all) 0.912 $\pm$ 0.104 0.890 $\pm$ 0.036 0.048
TabDDPM (10 discr. feat.) 0.998 $\pm$ 0.003 0.893 $\pm$ 0.044 0.016
TabDDPM (5 discr. feat.) 0.973 $\pm$ 0.068 0.896 $\pm$ 0.040 0.011
TabDDPM (2 discr. feat.) 0.968 $\pm$ 0.065 0.909 $\pm$ 0.039 0.005
VAE (baseline) 0.999 $\pm$ 0.005 0.360 $\pm$ 0.165 0.0
Table 15 shows results for Predictive Faithfulness (PF), Ground-truth Faithfulness (GF), and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency and variance, and using common motifs.
Table 15: Table showing results for Predictive Faithfulness (PF), Ground-truth Faithfulness (GF) and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency and variance using common motifs.
Approach PF GF Node feat.(CS)
Freq. $\uparrow$ $uparrow Var.\uparrow Freq. \uparrow$ Var. $\uparrow$
DBLP
DiTabDDPM (50 feat.) 0.875 $\pm$ 0.102 0.999 $\pm$ 0.0 0.723 $\pm$ 0.047 0.762 $\pm$ 0.037 0.233 0.229
DiTabDDPM (8 discr. feat.) 0.820 $\pm$ 0.091 0.882 $\pm$ 0.095 0.710 $\pm$ 0.065 0.668 $\pm$ 0.049 0.355 0.353
DiTabDDPM (4 discr. feat.) 0.749 $\pm$ 0.129 0.874 $\pm$ 0.078 0.688 $\pm$ 0.048 0.680 $\pm$ 0.074 0.387 0.386
DiTabDDPM (2 discr. feat.) 0.957 $\pm$ 0.065 0.965 $\pm$ 0.058 0.671 $\pm$ 0.058 0.712 $\pm$ 0.048 0.384 0.397
TabDDPM (50 feat.) 0.998 $\pm$ 0.001 0.996 $\pm$ 0.002 0.716 $\pm$ 0.046 0.729 $\pm$ 0.036 0.194 0.196
TabDDPM (8 discr. feat.) 0.986 $\pm$ 0.017 0.988 $\pm$ 0.023 0.709 $\pm$ 0.060 0.666 $\pm$ 0.052 0.221 0.221
TabDDPM (4 discr. feat.) 0.963 $\pm$ 0.038 0.963 $\pm$ 0.053 0.617 $\pm$ 0.057 0.635 $\pm$ 0.092 0.310 0.232
TabDDPM (2 discr. feat.) 0.974 $\pm$ 0.037 0.949 $\pm$ 0.071 0.647 $\pm$ 0.056 0.637 $\pm$ 0.062 0.235 0.235
VAE (baseline) 0.682 $\pm$ 0.048 0.601 $\pm$ 0.028 0.558 $\pm$ 0.066 0.612 $\pm$ 0.087 0.001 0.0
IMDB
DiTabDDPM (3066 feat., all) 0.979 $\pm$ 0.035 0.979 $\pm$ 0.035 0.915 $\pm$ 0.033 0.915 $\pm$ 0.033 0.029 0.029
DiTabDDPM (10 discr. feat.) 0.986 $\pm$ 0.034 0.983 $\pm$ 0.043 0.932 $\pm$ 0.022 0.902 $\pm$ 0.017 0.087 0.083
DiTabDDPM (5 discr. feat.) 0.972 $\pm$ 0.056 0.934 $\pm$ 0.098 0.923 $\pm$ 0.016 0.885 $\pm$ 0.022 0.072 0.069
DiTabDDPM (2 discr. feat.) 0.932 $\pm$ 0.096 0.980 $\pm$ 0.048 0.894 $\pm$ 0.022 0.905 $\pm$ 0.035 0.055 0.040
TabDDPM (3066 feat., all) 0.972 $\pm$ 0.065 0.972 $\pm$ 0.065 0.905 $\pm$ 0.035 0.905 $\pm$ 0.035 0.048 0.048
TabDDPM (10 discr. feat.) 0.953 $\pm$ 0.075 0.956 $\pm$ 0.085 0.894 $\pm$ 0.035 0.888 $\pm$ 0.025 0.017 0.016
TabDDPM (5 discr. feat.) 0.927 $\pm$ 0.083 0.991 $\pm$ 0.023 0.900 $\pm$ 0.036 0.914 $\pm$ 0.031 0.012 0.011
TabDDPM (2 discr. feat.) 0.980 $\pm$ 0.050 0.944 $\pm$ 0.078 0.904 $\pm$ 0.037 0.902 $\pm$ 0.030 0.008 0.005
VAE (baseline) 0.999 $\pm$ 0.0 0.999 $\pm$ 0.0 0.599 $\pm$ 0.0 0.599 $\pm$ 0.0 0.0 0.0
Experiments that consider dependence of node features
Table 16 shows results for Predictive Faithfulness (PF), Ground-truth Faithfulness (GF), and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency and variance.
Table 16: Table showing results for Predictive Faithfulness (PF) and Ground-truth Faithfulness (GF) and cosine similarity (Node Feat. (CS)) for feature selection experiments using frequency and variance.
Approach PF GF Node feat.(CS)
Freq. $\uparrow$ $uparrow Var.\uparrow Freq. \uparrow$ Var. $\uparrow$
DBLP
DiTabDDPM (50 feat.) 0.999 $\pm$ 0.0 1.0 $\pm$ 0.0 0.775 $\pm$ 0.047 0.905 $\pm$ 0.021 0.220 0.219
DiTabDDPM (8 discr. feat.) 0.952 $\pm$ 0.060 0.945 $\pm$ 0.061 0.800 $\pm$ 0.062 0.857 $\pm$ 0.053 0.298 0.302
DiTabDDPM (4 discr. feat.) 0.849 $\pm$ 0.056 0.816 $\pm$ 0.051 0.855 $\pm$ 0.057 0.825 $\pm$ 0.043 0.308 0.315
DiTabDDPM (2 discr. feat.) 0.898 $\pm$ 0.069 0.849 $\pm$ 0.081 0.837 $\pm$ 0.053 0.840 $\pm$ 0.033 0.279 0.272
TabDDPM (50 feat.) 0.999 $\pm$ 0.0 0.999 $\pm$ 0.0 0.877 $\pm$ 0.039 0.862 $\pm$ 0.047 0.227 0.226
TabDDPM (8 discr. feat.) 0.947 $\pm$ 0.081 0.953 $\pm$ 0.050 0.820 $\pm$ 0.066 0.825 $\pm$ 0.075 0.304 0.304
TabDDPM (4 discr. feat.) 0.835 $\pm$ 0.082 0.898 $\pm$ 0.047 0.812 $\pm$ 0.039 0.872 $\pm$ 0.068 0.311 0.311
TabDDPM (2 discr. feat.) 0.864 $\pm$ 0.058 0.875 $\pm$ 0.083 0.862 $\pm$ 0.037 0.862 $\pm$ 0.055 0.282 0.280
VAE (baseline) 0.664 $\pm$ 0.048 0.675 $\pm$ 0.049 0.717 $\pm$ 0.044 0.715 $\pm$ 0.032 0.0 0.0
IMDB
DiTabDDPM (3066 feat., all) 0.971 $\pm$ 0.067 0.971 $\pm$ 0.067 0.903 $\pm$ 0.045 0.903 $\pm$ 0.045 0.029 0.029
DiTabDDPM (10 discr. feat.) 0.974 $\pm$ 0.032 0.986 $\pm$ 0.039 0.926 $\pm$ 0.041 0.923 $\pm$ 0.047 0.088 0.085
DiTabDDPM (5 discr. feat.) 0.893 $\pm$ 0.105 0.878 $\pm$ 0.101 0.896 $\pm$ 0.045 0.919 $\pm$ 0.026 0.075 0.071
DiTabDDPM (2 discr. feat.) 0.918 $\pm$ 0.101 0.981 $\pm$ 0.048 0.906 $\pm$ 0.055 0.920 $\pm$ 0.059 0.061 0.044
TabDDPM (3066 feat., all) 0.979 $\pm$ 0.040 0.979 $\pm$ 0.040 0.919 $\pm$ 0.037 0.919 $\pm$ 0.037 0.048 0.048
TabDDPM (10 discr. feat.) 0.952 $\pm$ 0.060 0.978 $\pm$ 0.043 0.933 $\pm$ 0.044 0.883 $\pm$ 0.047 0.089 0.084
TabDDPM (5 discr. feat.) 0.971 $\pm$ 0.059 0.964 $\pm$ 0.070 0.920 $\pm$ 0.045 0.903 $\pm$ 0.048 0.077 0.075
TabDDPM (2 discr. feat.) 0.948 $\pm$ 0.063 0.953 $\pm$ 0.088 0.936 $\pm$ 0.043 0.903 $\pm$ 0.052 0.008 0.058
VAE (baseline) 0.999 $\pm$ 0.005 0.999 $\pm$ 0.005 0.36 $\pm$ 0.165 0.36 $\pm$ 0.165 0.0 0.0
Appendix K Hyperparameters
DiGNNExplainer
For GNN training, the hyperparameters for GraphSAGE include a learning rate of 0.005 and a weight decay of 0.001. The hyperparameters for HAN include a dropout of 0.6, a learning rate of 0.005, and a weight decay of 0.001. The GCNs use learning rate of 0.0005, weight decay of 5e-4 and dropout of 0.3. For graph generation using Diffusion Models, we use a batch size of 32, a learning rate of 0.0002, a weight decay of 1e-12, and diffusion steps of 5000. For node feature generation using discrete diffusion, we use a batch size of 32, a learning rate of 0.05, a weight decay of 1e-2, and diffusion steps of 25,000. The hyperparameters used for training in TabDDPM include a batch size of 1024, a learning rate of 0.002, a weight decay of 1e-4, and diffusion steps of 1000.
D4Explainer Baseline
Table 17 shows the hyperparameter settings that we used to run model-level explanations using D4Explainer for each of our datasets. The hyperparameters are chosen to support our GPU memory. We only train for 100 epochs since the training loss does not improve significantly after a few epochs. For a large dataset such as DBLP, the reduced number of training epochs is mainly due to the high computational time required without significant performance improvement.
Table 17: Model Parameters for D4Explainer baseline.
Dataset num hidden num layers batch size epochs
(in PPGN)
DBLP 64 6 8 20
IMDB 64 6 8 100
MUTAG 128 8 16 100
BA-Shapes 8 4 1 100
Tree-Cycle 8 4 16 100
Tree-Grids 8 4 16 100
BA-3Motif 128 6 16 100
Appendix L Implementation
The code for implementation is written in Python 3.10.6. We utilized the following Python libraries—Pandas for data manipulation, matplotlib and seaborn for visualizing plots, Networkx for plotting graphs, PyTorch Geometric for GNN Training, sklearn for sampling node features and comparing feature distributions, Forest Fire Sampler for sampling subgraphs, Deep Graph Library (DGL) for creating metagraphs, Python-louvain for community detection, and TransE implementation of pykeen library for generating node embeddings.
The experiments were conducted on a machine with Ubuntu 22.04.2 LTS operating system, AMD Ryzen 9 CPU, and a NVIDIA GeForce RTX 3070 GPU with 8GB of GPU memory.

### v2

@@ Appendix D Validity Analysis
We observe that a higher number of graphs generated using DiGNNExplainer are valid (i.e., consistent with the metagraph of the dataset) for DBLP compared to IMDB (Table 8). For homogeneous datasets, the number of connected graphs is the same as valid graphs, as homogeneous graphs do not require validation using a metagraph.
Table 8. Analysis of the validity of generated graphs according to the metagraph.
@@
Appendix E Community Detection
We use communities or motifs to evaluate the ground-truth faithfulness of the explanation graphs (see Section 4.1), following previous work (Ying et al., 2019; Yuan et al., 2020). Table 9 (right) shows example motifs.
For real-world graphs (DBLP, IMDB, MUTAG), we employ Louvain community detection [^6] https://python-louvain.readthedocs.io/en/latest/api.html (Blondel et al., 2008) which identifies the largest substructures in input graphs. Unlike existing techniques (Subramonian, 2021; Yu and Gao, 2022) that often rely on blackbox models for motif detection, our method is more similar to GraphXAI (Agarwal et al., 2023), which detects ground-truth substructures in the real-world dataset MUTAG via substructure matching. For synthetic graphs (BA-3Motif, BA-SHAPES, Tree-Cycle, and Tree-Grids), we generate motifs using ShapeGGen (Agarwal et al., 2023).
We use class-specific motifs when motifs can be associated with a prediction class, and common motifs otherwise. For the graph classification datasets (MUTAG, BA-3MOTIF), each graph belongs to a single class, so we construct class-specific motif sets. Similarly, for the real-world node classification datasets DBLP and IMDB, we obtain class-specific motifs. In DBLP, most (104 out of 120) of the sampled graphs using ForestFireSampler contain author nodes from a single class, in IMDB, 40 out of 143 graphs contain movie nodes from a single class. If a sampled graph has nodes of different classes (e.g., containing author nodes from classes 1 and 3), we do not use it for motif detection, because the obtained motifs might be common across classes. In this way, we detect motifs that help to distinguish different classes.
In contrast, for the synthetic node classification datasets BA-Shapes, Tree-Cycle, and Tree-Grids, the task is to predict node labels and every graph contains nodes of all labels. For example, in BA-Shapes, a house motif has all four node labels. As a result, motifs are naturally shared across labels/classes, and we evaluate faithfulness using a single set of common motifs. Table 10 shows the statistics for class-specific and common motifs for our datasets.
Table 9. Explanation graphs for Class 1
@@ Appendix E Community Detection
Table 10. Motif statistics. Top: For datasets with class-specific motifs, we report, for each class, the number of detected motifs that occur and do not occur in the explanation graphs. Bottom: For datasets with common motifs, we report the total number of motifs and how many motifs appear in the best explanation graph of every class, in at least $\frac{1}{2}$ of the classes, and in none of the classes.
Dataset Class 0 Class 1 Class 2 Class 3
occur not occur occur not occur occur not occur occur not occur
DBLP 8 2 8 2 10 0 8 2
IMDB 9 1 9 1 10 0
MUTAG 30 14 3 11
BA-3Motif 7 3 2 8 10 0
Dataset No. of motifs No. of motifs No. of motifs No. of motifs
(every class) ($\frac{1}{2}$ of the classes) (none of the classes)
BA-Shapes 30 14 26 4
Tree-Cycle 30 15 28 2
Tree-Grids 30 28 30 0
Table 11. Feature selection by frequency
@@
DiTabDDPM (50 feat.) 0.233 0.902 $\pm$ 0.102 0.867 $\pm$ 0.046
DiTabDDPM (8 discr. feat.) 0.355 0.797 $\pm$ 0.062 0.852 $\pm$ 0.054
DiTabDDPM (4 discr. feat.) 0.387 0.770 $\pm$ 0.121 0.820 $\pm$ 0.055
DiTabDDPM (2 discr. feat.) 0.384 0.972 $\pm$ 0.041 0.845 $\pm$ 0.052
TabDDPM (50 feat.) 0.194 0.998 $\pm$ 0.001 0.842 $\pm$ 0.054
TabDDPM (8 discr. feat.) 0.221 0.976 $\pm$ 0.034 0.852 $\pm$ 0.052
TabDDPM (4 discr. feat.) 0.310 0.916 $\pm$ 0.072 0.857 $\pm$ 0.057
TabDDPM (2 discr. feat.) 0.235 0.969 $\pm$ 0.038 0.822 $\pm$ 0.042
VAE (baseline) 0.0 0.664 $\pm$ 0.048 0.717 $\pm$ 0.044
@@
DiTabDDPM (3066 feat., all) 0.029 0.955 $\pm$ 0.084 0.890 $\pm$ 0.029
DiTabDDPM (10 discr. feat.) 0.083 0.977 $\pm$ 0.066 0.933 $\pm$ 0.036
DiTabDDPM (5 discr. feat.) 0.069 0.970 $\pm$ 0.065 0.900 $\pm$ 0.036
DiTabDDPM (2 discr. feat.) 0.040 0.984 $\pm$ 0.039 0.913 $\pm$ 0.047
TabDDPM (3066 feat., all) 0.048 0.912 $\pm$ 0.104 0.890 $\pm$ 0.036
TabDDPM (10 discr. feat.) 0.016 0.998 $\pm$ 0.003 0.893 $\pm$ 0.044
TabDDPM (5 discr. feat.) 0.011 0.973 $\pm$ 0.068 0.896 $\pm$ 0.040
TabDDPM (2 discr. feat.) 0.005 0.968 $\pm$ 0.065 0.909 $\pm$ 0.039
VAE (baseline) 0.0 0.999 $\pm$ 0.005 0.360 $\pm$ 0.165
Appendix F Case Study—Explanations for the same class
Table 9 shows a qualitative analysis of the explanation graphs obtained for the same class (class 1) for all datasets and all approaches, along with example motifs for each dataset.
Appendix G Feature Selection by Frequency
@@
Table 11 reports predictive faithfulness (PF), ground-truth faithfulness (GF), and cosine similarity (Node Feat. (CS)) for frequency-based feature selection.
Appendix H Hyperparameters
For GNN training, we use the following settings. GraphSAGE is trained with a learning rate of 0.005 and a weight decay of 0.001. HAN is trained with a dropout of 0.6, a learning rate of 0.005, and a weight decay of 0.001. For GCN training, on MUTAG, we use a dropout of 0.3 and a learning rate of 0.0005, whereas for synthetic datasets we use 0.5 and 0.005, respectively.
For graph generation with diffusion models, we use a batch size of 256, a learning rate of 0.0002, a weight decay of 1e-12, and 5,000 diffusion steps. For DiTabDDPM, we use a batch size of 32, a learning rate of 0.05, a weight decay of 0.001, and 25,000 diffusion steps. For TabDDPM, we use the same hyperparameters as in the original implementation.
Appendix I Implementation
All code is implemented in Python 3.10.6. We utilize Pandas for data manipulation, Matplotlib and Seaborn for visualization, NetworkX for graph plotting, PyTorch Geometric for GNN training, scikit-learn for sampling node features and comparing feature distributions, the Forest Fire Sampler for sampling subgraphs, the Deep Graph Library (DGL) for creating metagraphs, and Python-louvain for community detection.
The experiments were conducted on a machine with Ubuntu 22.04.2 LTS, an AMD Ryzen 9 CPU, and an NVIDIA GeForce RTX 3070 GPU with 8GB of GPU memory.